OpenStudy (anonymous):

Multiply: √ 3(2 + √ 12)

OpenStudy (saifoo.khan):

\[\frac{\sqrt{245}}{2\sqrt5}\]? was that line like this?

OpenStudy (lilg132):

6 + 2 sqrt 3

OpenStudy (lilg132):

6 \[+2\sqrt{3}\]

OpenStudy (lilg132):

\[6+2\sqrt{3}\]